# Explainable Multi-Modal Data Exploration in Natural Language via LLM Agent

### v1

@@
Abstract.
International enterprises, organizations, or hospitals collect large amounts of multi-modal data stored in databases, text documents, images, and videos. While there has been recent progress in the separate fields of multi-modal data exploration as well as in database systems that automatically translate natural language questions to database query languages, the research challenge of querying database systems combined with other unstructured modalities such as images in natural language is widely unexplored.
@@ Abstract.
In this paper, we propose XMODE  [^1] The source code, data, and/or other artifacts have been made available at https://github.com/yizhang-unifr/XMODE. - a system that enables explainable, multi-modal data exploration in natural language. Our approach is based on the following research contributions: (1) Our system is inspired by a real-world use case that enables users to explore multi-modal information systems. (2) XMODE leverages a LLM-based agentic AI framework to decompose a natural language question into subtasks such as text-to-SQL generation and image analysis. (3) Experimental results on multi-modal datasets over relational data and images demonstrate that our system outperforms state-of-the-art multi-modal exploration systems, excelling not only in accuracy but also in various performance metrics such as query latency, API costs, planning efficiency, and explanation quality, thanks to the more effective utilization of the reasoning capabilities of LLMs.
1. Introduction
Consider a hospital in the near future in which doctors, nurses, and data scientists naturally access digital patient data. This data includes electronic health records (EHR), usually stored in relational databases (sivasubramaniamsm3, ), but also multimedia data such as medical images from CT scans or X-rays and the corresponding reports written by medical experts (unstructured data). Each participant seeks to interactively query all these datasets in natural language. Different participants also have different skill sets and exploration goals. Additionally, given the application domain, each user wants to understand exactly how the system evaluates their queries. A system that supports such a scenario would unlock a plethora of applications, from this medical example to queries over shared scientific databases (also containing structured data, text, images, and videos), queries over public datasets, and more. However, building such a system presents significant research challenges in understanding user intent, which often relies on complex queries, querying multimedia databases, and ensuring explainability.
(a) Medical Data Exploration
(b) Art Work Data Exploration
Figure 1. Example workflows of multi-modal data exploration in natural language over heterogeneous data sources. A complex natural language question is decomposed into sub-questions to better enable answer explainability. Each sub-question is designated to a particular task (such as text-to-SQL translation or image analysis). These tasks may be expanded to utilize various tools and machine learning models to address specific downstream requirements necessary for answering a user’s natural language question.
To understand these challenges, a concrete scenario of multi-modal exploration involving a relational database, text documents, and images is outlined here.. Assume that a user asks the following question in natural language: Show me the progression of cancer lesions over the last 12 months of patients with lung cancer who are smokers. (see the upper part of Figure 1 b). This seemingly straightforward query encapsulates several fundamental challenges in multifaceted data exploration. First, it requires the decomposition of a natural language query into semantically precise sub-queries, each targeting diverse data modalities while preserving the original intent. Critical to this process is optimizing the workflow sequence - determining which queries should be executed first to minimize computational overhead and maximize efficiency. For instance, filtering patients through structured database queries before retrieving and analyzing medical images significantly reduces the computational burden compared to analyzing all available images first. In our example in Figure 1, natural language $NL1$ is a text-to-SQL task to query the relational database for the name and age of patients diagnosed with lung cancer. The result is then used for $NL2$ an image analysis task - looking for cancer lesions in those patients’ images. Finally, $NL3$ a visualization task - shows the cancer progression for each patient. This workflow sequence is deliberately optimized: starting with structured data filtering before proceeding to more computationally intensive image analysis tasks. The complexity compounds when considering the temporal aspect of disease progression, which necessitates careful alignment of data across different modalities and timestamps. Furthermore, in healthcare settings, result verification and transparency are paramount. Users must be able to trace back any conclusions to the source data, understand how intermediate results were derived, and verify the accuracy of each analytical step. This necessitates a workflow where users can validate intermediate results before proceeding to subsequent analysis stages, with the ability to refine queries if the results don’t align with their clinical expectations.
Now imagine a museum or art gallery in the near future, where curators, researchers, and data scientists can naturally access and explore digital art collections. This data includes structured information about paintings, such as artist, title, medium, and subject matter, which is typically stored in a relational database. The collection also includes unstructured data, such as the full text of art critiques and descriptions, as well as digital images of the artworks. Similar to the previous use case for medical data exploration, also the use case for artwork exploration is multi-modal and requires analysis of heterogeneous data (e.g. tabular and image) as shown in Figure 1(b).
@@ 1. Introduction
  • Heterogeneous data exploration: How can we design a system that accurately interprets user queries in natural language for exploring heterogeneous data sources with high accuracy?
  • Orchestrating multiple expert models and tools for data exploration: How can we automatically break down a user question into sub-questions that can later be organized into a workflow plan? How do we delegate these tasks to the appropriate expert models from the available toolbox, considering dependencies and the potential for parallel execution?
  • Explainability: How can we design a system that facilitates multi-modal exploration, allowing end users to trace conclusions back to their source data, comprehend how intermediate results were generated, and identify situations where questions remain unanswered due to missing data?
Existing works on multi-modal data exploration in natural language follow mainly two paradigms (1) multiple modalities are embedded in a single query language, e.g., NeuralSQL (bae2024ehrxqa, ) embeds visual QA functions directly in SQL; (2) agentic workflows, in which different tools (e.g., relational database operators, vision model) are intelligently combined to answer a user-question such as Caesura (urban2023caesura, ).
In this paper, we propose XMODE - a multi-modal data exploration system that uses an Large language Model based agentic framework to tackle these challenges. The basic idea is to first decompose a complex natural language question into simpler sub-questions. Each question is then translated into a workflow of specific tasks. By applying smart planning, our approach can reason about which task in the workflow fails and thus re-plan that specific task rather than restarting the complete workflow. The advantage of our approach compared to similar systems such as Caesura (urban2023caesura, ) is that it enables parallel task execution through the construction of a directed acyclic task graph and requires a lower number of tokens from prompt engineering resulting in more efficient query execution times and API calling costs. The main contributions of our paper are as follows:
  • Higher accuracy: XMODE is based on an agentic AI framework that shows higher accuracy for exploring multi-modal data than traditional work due to the smart orchestration of different tasks of the data exploration pipeline.
  • Improved performance: XMODE demonstrates performance improvements compared to state-of-the-art through parallelism, reasoning and smart re-planning.
  • Better explainability: XMODE enhances explainability by enabling a user to inspect the decisions and reasoning at each step that led to the final output, tracing back through the results of all previous steps.
  • Generalizablilty: XMODE is designed and evaluated in a zero-shot setting, demonstrating its ability to perform complex tasks without relying on In-Context Learning (ICL), thereby improving both adaptability and accessibility.
2. Related Work
@@
The research field of text-to-SQL systems has seen tremendous progress over the last few years (Floratou2024, ; pourreza2024din, ) due to advances in large language models. Original success can be attributed to rather simplistic datasets consisting of databases with only several tables as in Spider (yu2018spider, ). Especially the introduction of new benchmarks such as ScienceBenchmark (zhang2023sciencebenchmark, ) or BIRD (li2024can, ) has further pushed these limits of these systems. Most of the research efforts have been restricted to querying databases in English apart from a few exceptions such as Statbot.Swiss (noor2024, ).
Explainability.
Explainability aims to provide a deeper understanding of how machine learning models make predictions by illuminating the decision-making processes within these models. It strives to offer transparency, enabling stakeholders to comprehend, trust, and effectively manage the outcomes produced by these models (lundberg2017unified, ; kim2018interpretability, ). Although there has been recent progress in artificial intelligence in general, for the task of data exploration in natural language, explainability is an open issue. Recently, in the multi-agent collaboration framework (wang2023towards, ), explainability has been designed to mimic human-like top-down reasoning by utilizing the extensive knowledge of Large Language Models (LLMs). For the task of text-to-SQL, explainability is basically an unexplored research topic with the exception of back-translating automatically generated SQL statements to natural language (bandyopadhyay2020natural, ; von2022improving, ; zhang2023sciencebenchmark, ). However, back translation is often not enough to fully explain how a system comes up with an answer and how to interpret the results.
Multi-modal systems.
@@
Video Database Management Systems (VDBMSs) support efficient and complex queries over video data, but are often restricted to videos only (e.g., (zhang2023equi, ; DBLP:journals/pvldb/KangBZ19, ; DBLP:conf/deem/KakkarCCXVDPBS023, )). ThalamusDB (jo2024, ) enables queries over multi-modal data but requires SQL as input, with explicit identification of the predicates that should be applied to an attribute corresponding to video or audio data. Similarly, MindsDB [^2] https://docs.mindsdb.com and VIVA (DBLP:conf/cidr/KangRBKZ22, ) require that users write SQL and manually combine data from relational tables and models. Vision-language models provide textual descriptions of video data (zhang2024vision, ), but are not designed to support precise, structured queries.
Most closely related to our approach are CAESURA (urbanB24, ), which supports natural language queries over multi-modal data lakes, and PALIMPZEST (liu2024declarative, ), which enables optimizing AI workload. The key distinction of our system, XMODE, is its focus on efficiently orchestrating various model calls and their dependencies. This approach not only improves latency and cost but also enhances accuracy by minimizing interference from the outputs of intermediate function calls.
Moreover, the related systems enable multi-modal queries across structured and unstructured data with a focus on query planning. However, these systems do not address enhancing the accuracy and explainability of the underlying model for natural language data exploration tasks. Explainability and answer justification are crucial in domains like medical data science, where medical device regulations mandate systems to provide detailed explanations of how specific results are obtained, ensuring that no potentially fatal medical treatment is recommended.
Figure 2. XMODE system architecture.
3. System Design
We now describe the design of our system called XMODE, which enables explainable multi-modal data exploration in natural language.
3.1. System Architecture of XMODE
Figure 3. XMODE system architecture in ArtWork  (urban2023caesura, ) with an example of processing a multi-modal query. The query is automatically decomposed into various components such as text2SQL, and image analysis which can be inspected by the user for explainability.
Figure 4. XMODE system architecture in EHRXQA  (bae2024ehrxqa, ) with an example of processing a multi-modal query. The query is automatically decomposed into various components which can be inspected by the user for explainability.
The architecture of our system, XMODE, is illustrated in Figures 2. We describe the five primary components of XMODE using an example query applied to artwork data, which includes relational tables and images: Plot the number of paintings that depict war for each century. The system’s operation is depicted in Figure  3.
@@ 3.1. System Architecture of XMODE
XMODE is an agentic system  (kapoor2024ai, ) driven by a llm-based dynamic planner pattern  (kim2023llm, ) equipped with a comprehensive toolkit containing all the necessary models to decompose a user’s question, such as a multi-modal natural language question, into a workflow (i.e., a graph of sub-questions). The workflow is represented as a Directed Acyclic Graph (DAG), of which each node corresponds to a simple sub-question with a specific tool assigned by the planner. The planner determines sub-tasks that can be executed in parallel and it manages their dependencies. XMODE is designed to be adaptable and to allow for dynamic debugging and plan modification (re-planning) if necessary, e.g., in case of failures during a text-to-SQL sub-task. As it shown in Figure 1, the design of XMODE incorporates multiple components:
  1. (1)
    Planning & Expert Model Allocation. The system analyzes the user question, then constructs a sequence of tasks to be executed considering their dependency. It determines the required expert models from an available toolkit for each task, as well as their input arguments and their inter-dependencies to synthesize them as a workflow. To do so, it employs the power of reasoning capability of LLMs. The output of this stage is a workflow in the form of a DAG that formalizes task dependencies. As we can see in Figure 3, the original natural language question is split into four tasks $t_{1}$ to $t_{4}$, namely text2SQL, image_analysis, data_preparation and data_plotting. We leverage LLMs’ reasoning ability to generate a workflow from natural language questions by providing detailed specifications of each expert model available in the toolkit.
  2. (2)
    Execution and Self-Debugging. The system executes tasks according to a generated workflow by calling allocated expert models from a toolkit. The system employs a state object that stores all the intermediate interactions during the execution of a workflow. The independent tasks are executed concurrently and after completing each task, the outcomes are passed on as input to the tasks that rely on them according to the workflow. Each expert model has an inner self-debugging component to handle errors that can occur during its execution. As we can see in the middle of Figure 3, XMODE provides reasoning in natural language for each task which can easily be understood by humans.
  3. (3)
    Decision Making. In this part, XMODE synthesizes results from a state object and inspects them for a final decision. If the task results are sufficient to fulfill a user request, it will prepare the final results to respond to the user, otherwise it will request the planning component to re-plan a new workflow by providing the intermediate results and reasons for re-planning. This process repeats until the decision making components are satisfied with the final outcome to present to the user or the maximum loop limit is reached. This component benefits from the reasoning capability of LLMs in the decision making process.
  4. (4)
    Expert Models & Tools. This component contains expert models such as machine learning models that perform specific downstream tasks such as text-to-SQL, image analysis, and text analysis. It also contains particular tools such as data formatting and plotting tools. Taking into account various use cases, the toolkit section of XMODE provides access to these models and tools. Each expert model or tool should include a description and argument specifications and they will be available to the planning module.
  5. (5)
    Data Lake. A repository containing structured and unstructured data such as tabular data, images, and text. Each model expert and tool has direct access to the repository to conduct the assigned task.
Our current XMODE implementation offers a range of features, including query debugging, query re-planning, optimization and explainability to better understand how a natural language question is decomposed into multiple sub-tasks. Each feature of our system is available at varying levels of complexity.
4. Experiments
In this section, we evaluate the performance of our system XMODE. In particular, we want to address the following research questions:
  • How well does XMODE tackle multi-modal natural language questions on two different datasets consisting of tabular data, and images?
  • How does the system perform compared to state-of-the-art systems such as CAESURA (urbanB24, ) and NeuralSQL (bae2024ehrxqa, )?
  • Which explanations does the system provide to justify the answers?
4.1. Experimental Setup
4.1.1. Datasets
For our experiments, we used two different datasets, namely information about artwork as well as electronic health records.

### v2

# Multi-Modal Data Exploration via Language Agents

@@
Abstract
@@
International enterprises, organizations, and hospitals collect large amounts of multi-modal data stored in databases, text documents, images, and videos. While there has been recent progress in the separate fields of multi-modal data exploration as well as in database systems that automatically translate natural language questions to database query languages, the research challenge of querying both structured databases and unstructured modalities (e.g., texts, images) in natural language remains largely unexplored. In this paper, we propose M2EX  [^1] Data and code repository are available at https://github.com/yizhang-unifr/M2EX—a system that enables multi-modal data exploration via language agents. Our approach is based on the following research contributions: (1) Our system is inspired by a real-world use case that enables users to explore multi-modal information systems. (2) M2EX leverages an LLM-based agentic AI framework to decompose a natural language question into subtasks such as text-to-SQL generation and image analysis and to orchestrate modality-specific experts in an efficient query plan. (3) Experimental results on multi-modal datasets, encompassing relational data, text, and images, demonstrate that our system outperforms state-of-the-art multi-modal exploration systems, excelling in both accuracy and various performance metrics, including query latency, API costs, and planning efficiency, thanks to the more effective utilization of the reasoning capabilities of LLMs.
Multi-Modal Data Exploration via Language Agents
Farhad Nooralahzadeh, Yi Zhang, Jonathan Fürst, Kurt Stockinger Zurich University of Applied Sciences, Switzerland {farhad.nooralahzadeh, yi.zhang, jonathan.fuerst, kurt.stockinger}@zhaw.ch
1 Introduction
The rapid expansion of multi-modal data; spanning structured tables, text, images, and video; has created an urgent need for flexible, scalable systems for complex data exploration. In fields like healthcare, users often query across EHRs, medical images, and clinical notes using natural language. However, current systems struggle with integrating modalities, capturing user intent, and optimizing execution workflows, limiting their real-world utility. Traditional solutions focus on single-modality tasks such as text-to-SQL (Sivasubramaniam et al., 2024; Nooralahzadeh et al., 2024; Pourreza and Rafiei, 2024), visual question answering (Li et al., 2023a; Ko et al., 2023; Du et al., 2023), or domain-specific QA (Dong et al., 2024; Liu et al., 2024b), often relying on rigid pipelines or handcrafted logic. While effective in narrow settings, they lack the flexibility to handle heterogeneous data or dynamic analytical goals.
Recent large language models (LLMs) and vision-language models (VLLMs) offer broader generalization but remain limited in real-world multimodal use. Techniques like retrieval-augmented generation (RAG) improve grounding but often fail at structured reasoning, long-term context, and precise tool use; especially in domains that require deep alignment across modalities and step-wise execution.
Figure 1: (Left): Example workflows of multi-modal data exploration in natural language over heterogeneous data sources. (Right): M2EX system architecture.
Efforts to inspire LLMs with agentic capabilities – such as ReAct (Yao et al., 2023), tool invocation (Yang et al., 2023; Schick et al., 2023), or workflow automation (Liu et al., 2024a; Urban and Binnig, 2024) – have further exposed systemic challenges.
Existing frameworks frequently adopt rigid, sequential decision-making processes, incurring computational overhead and limiting scalability. Evaluations of these systems are often conducted on in-house datasets, lacking rigorous benchmarking against ground-truth metrics or real-world multi-modal contexts. Moreover, many approaches enforce fixed task-planning hierarchies or routing mechanisms, stifling adaptability and reusability across diverse applications. This “one-size-fits-all” mentality contrasts starkly with the need for modular, composable agents capable of dynamically integrating domain-specific tools, retaining contextual memory, and self-optimizing workflows.
To understand these challenges, a concrete scenario of multi-modal exploration involving a relational database, text documents, and images is outlined here. A seemingly straightforward query like Show me the progression of cancer lesions over the last 12 months of patients with lung cancer who are smokers (see Figure 1, Left) requires multi-modal integration, posing challenges in decomposition and optimization. Critical to this process is optimizing the workflow sequence, i.e., determining which queries should be executed first to minimize computational overhead and maximize efficiency.
In this work, we propose a novel framework for multi-modal data exploration that bridges these gaps through LLM-based agents designed for extensibility, precision, and cross-domain generalization. Our approach combines a “Swiss army knife” philosophy — enabling reusable, adaptable modules for tasks like semantic parsing, cross-modal retrieval, and structured data operations — with a principled evaluation strategy spanning diverse benchmark datasets. By decoupling task planning from execution and incorporating feedback-driven memory, our system supports iterative exploration while mitigating the pitfalls of shallow evaluation and fixed workflows. We demonstrate its efficacy across text, visual, tabular, and hybrid data domains, underscoring the potential of agentic LLMs to unify multi-modal analysis in a scalable, user-centric paradigm.
@@ 1 Introduction
  • Heterogeneous data understanding: How can we accurately interpret natural language queries over diverse data types such as text, tables, and images?
  • Workflow orchestration: How can we decompose a complex query into sub-tasks, organize them into an executable workflow, and delegate each to the right model or tool; respecting dependencies and enabling parallelism?
  • Explainability: How can we provide users with traceable, transparent results, showing how answers were derived, what data contributed, and where uncertainty remains?
In this paper, we propose M2EX—a multi-modal data exploration system that uses a LLM-based agentic framework to tackle these challenges. The basic idea is to first decompose a complex natural language question into simpler sub-questions. Each sub-question is then translated into a workflow of specific tasks. By applying smart planning, our approach can reason about which task in the workflow fails and thus re-plan that specific task rather than restarting the complete workflow. The advantage of our approach compared to similar systems such as CAESURA (Urban and Binnig, 2024) is that it enables parallel task execution through the construction of a directed acyclic task graph and requires a lower number of tokens from prompt engineering, resulting in more efficient query execution times and API calling costs.
The main contributions of our paper are as follows: (i) Unified DAG–first planning.The planner compiles a natural-language query directly into an execution directed-acyclic graph (DAG); independent subtasks therefore run in parallel without a second “physical-plan” stage. (ii) Self-debug & selective re-planning for speed.Each expert tool validates its own output once; if a fault persists, the agent rewires only the affected sub-graph. This cuts end-to-end latency by up to 51% and reduces token usage by 18% on the ArtWork benchmark. (iii) Zero-shot cross-domain generalisation.With a single prompt set and no in-context examples, M2EX attains up to 42% higher answer accuracy than CAESURA and NeuralSQL on ArtWork, RotoWire, and EHRXQA.
2 Related Work
@@
The research field of text-to-SQL systems has seen tremendous progress over the last few years (Floratou et al., 2024; Pourreza and Rafiei, 2024) due to advances in large language models. Original success can be attributed to rather simplistic datasets consisting of databases with only several tables, as in Spider (Yu et al., 2018). Especially the introduction of new benchmarks such as ScienceBenchmark (Zhang et al., 2024b), FootbalDB (Fürst et al., 2024), BIRD (Li et al., 2024b) or SM3 (Sivasubramaniam et al., 2024) has further pushed the limits of these systems. Most of the research efforts have been restricted to querying databases in English apart from a few exceptions such as Statbot.Swiss (Nooralahzadeh et al., 2024).
@@
Multi-modal systems. Video Database Management Systems (VDBMSs) support efficient and complex queries over video data, but are often restricted to videos only (e.g., Zhang et al., 2023; Kang et al., 2019; Kakkar et al., 2023). ThalamusDB (Jo and Trummer, 2024) enables queries over multi-modal data but requires SQL as input, with explicit identification of the predicates that should be applied to an attribute corresponding to video or audio data. Similarly, MindsDB [^2] https://docs.mindsdb.com and VIVA (Kang et al., 2022) require that users write SQL and manually combine data from relational tables and models. Vision-language models provide textual descriptions of video data (Zhang et al., 2024a), but are not designed to support precise, structured queries. Recent multi-modal systems such as MAGMA Doe et al. (2025), and LLaVA-Next Li et al. (2024a) extend vision-language reasoning via unified interfaces or tool-based controllers. However, these models are largely limited to vision-only pipelines and lack support for structured tool orchestration across modalities. In contrast, M2EX generalizes to diverse tool types—including text-to-SQL, Python plotting, and image-VQA—via explicit DAG planning and partial re-planning, enabling scalable and interpretable execution across multi-modal queries.
Closest to our work are CAESURA (Urban and Binnig, 2024), PALIMPZEST (Liu et al., 2024a), and MAT Gao et al. (2025), which address multi-modal querying and AI workload optimization. In contrast, M2EX focuses on efficient orchestration of model calls and dependencies, reducing latency and cost while improving accuracy by minimizing interference from intermediate outputs (Schick et al., 2023) [^3] CAESURA and MAT employ the ReAct agent framework, which leads to extended context tokens and increased latency..
While related systems emphasize query planning, they fall short in enhancing the accuracy and explainability of model outputs—critical needs in domains like medical data science, where regulatory standards require transparent and justifiable results.
3 Method and System Design
Problem statement.
Given a multi-modal query $q$, a data lake $D$, a tool catalogue $T$ with metadata $T_{\text{meta}}$, our goal is to produce a directed acyclic task graph $G=(V,E)$ and a final answer $a$ such that each node $v\!\in\!V$ is a (tool, args) pair, edges $E$ encode data dependencies, the execution of $G$ is valid w.r.t. $T_{\text{meta}}$, and $a$ maximizes task-level answer accuracy.
Proposed System:
@@ 3 Method and System Design
To address this problem, M2EX enables multi-modal data exploration via language agents. Its details are presented in Algorithm 1 and Figure 1 (right) (A fully annotated DAG and end-to-end use-case example appear in Figures 2 and 3). M2EX is an agentic system (Kapoor et al., 2024) driven by LLMCompiler (Kim et al., 2023), a dynamic planner pattern based on a Large Language Model, equipped with a comprehensive toolkit $\mathcal{T}$ containing all the necessary models to decompose a user’s request, such as a multi-modal natural language question, into a workflow (i.e., a graph of sub-questions). The workflow is represented as a Directed Acyclic Graph (DAG), where each node corresponds to a simple sub-task (or sub-question) with a specific tool assigned by the planner. While decoupling logical and physical plans can be suboptimal due to plan ambiguity and nonlinearity, unlike CAESURA, the planner determines sub-tasks that can be executed in parallel and manages their dependencies by leveraging an LLM to directly generate the execution plan from the query as a graph of function calls. M2EX is designed to be adaptable, allowing dynamic debugging and plan modification (re-planning) when necessary, for example, if a failure occurs during a text-to-SQL sub-task.
Algorithm 1 M2EX: Multi-Modal Data Exploration via Language Agents
1User query $q$, Agent Core $\mathcal{LLM}$, toolkit $\mathcal{T}$, Data Lake $D$, Pre-defined Prompts $\mathcal{P}$, Empty memory state $\mathcal{R}$
2Final answer $a$
3Stage 1: Planning & Expert Model Allocation
[…]
7Stage 2: Execution & Self-debugging
[…]
10for each batch $b_{k}\in\mathcal{B}$ do
11  Launch parallel execution:
12  for each subtask $s_{i}\in b_{k}$ do
13   $r_{i}\leftarrow\textsc{Execute}(s_{i},\mathcal{T},\mathcal{D})$ $\triangleright$ Invoke the assigned expert tool for $s_{i}$. Integrate $n$-time self-debugging to automatically detect and correct errors as needed. ($n=1$). If there is still an error, provide an error message as an output of execution.
14   $\mathcal{R}\leftarrow\mathcal{R}\cup\{r_{i}\}$
15  end for
16end for
17Stage 3: Decision Making
18Validate $\mathcal{R}$ via reflection $\triangleright$ Check that outputs are correct and executable; if not, trigger error feedback.
19if validation fails then
20  $G\leftarrow\textsc{Replan}(G,\mathcal{R},\mathcal{LLM},\ \mathcal{T}_{meta})$ $\triangleright$ Dynamically adjust the DAG (e.g., reallocate tasks or update tool parameters) based on error feedback using an agent core $\mathcal{LLM}$ (with a replanning prompt $\in\mathcal{P}$).
21  goto line 5 $\triangleright$ Restart execution with the updated plan.
22end if
[…]
24if $a$ is insufficient or uncertain then
25  $G\leftarrow\textsc{Replan}(G,\mathcal{R},\mathcal{LLM},\ \mathcal{T}_{meta})$ $\triangleright$ Dynamically adjust the DAG (e.g., reallocate tasks or update tool parameters) based on error feedback using an agent core $\mathcal{LLM}$ (with a replanning prompt $\in\mathcal{P}$).
26  goto line 5 $\triangleright$ Restart execution with the updated plan.
27end if
28return $a$
As shown in Algorithm 1 and Figure 1, the system is composed of the following key components: (1) User Query ($q$):a multi-modal natural language question posed by the user, which initiates the process of task decomposition and execution. (2) Agent Core ($\mathcal{LLM}$):the core reasoning engine that powers the dynamic planning, execution, and decision-making processes. The LLM is responsible for decomposing the user query into subtasks, managing dependencies, and synthesizing final results using diverse prompts $\mathcal{P}$. (3) Expert Models & Tools ( Toolkit ) ($\mathcal{T}$):a comprehensive collection of expert models and tools that are used for executing specific sub-tasks. The toolkit provides the necessary models for tasks such as text-to-SQL, text analysis, image analysis, data preparation, and data plotting. Each expert model or tool should include a description and argument specifications ($\mathcal{T}_{meta}$), and they will be available during the planning and re-planning stages. (4) Data Lake ($D$):a central repository that stores both structured and unstructured data, such as tabular data, images, and text. Each expert model and tool has direct access to the data lake to perform its assigned tasks. The data stored in the lake is utilized as input for various tasks, enabling the system to generate accurate results for the user’s query. (5) Pre-defined Prompts ($\mathcal{P}$):a collection of predefined prompts available to the LLM, which are used to guide the reasoning process during planning, execution, and decision-making (see details in Appendix B). (6) Memory State ($\mathcal{R}$):The initial memory state starts empty and captures all intermediate results and interactions throughout the workflow execution. The system tracks these intermediate results using an output object that stores the answer and reasoning at each node in the workflow. (7) Final Answer ($a$):The final answer is the output generated by the system after executing all the tasks and performing reasoning through the LLM. It consolidates all intermediate results and provides a comprehensive response to the user’s query. The final answer typically includes several components: a summary of the task or query result, detailed information about the outcome, the source of the data used, an inference indicating the success of the task, and any additional explanations or clarifications. This structured output ensures that the user receives not only the result but also the reasoning and context behind it. In Figure  2, we demonstrate the showcase of M2EX using an example query applied to the EHRXQA data, which includes relational tables and images: Was patient 18061894 prescribed acetaminophen, and did a chest x-ray show any technical assessments until 12/2103?
The system starts with the user query $q$ and processes it through several stages, as detailed below:
  1. (i)
    Planning & Expert Model Allocation. The system begins by analyzing the user query $q$ and decomposes it into a sequence of tasks. Using the agent core ($\mathcal{LLM}$), the system identifies the required expert models and tools from the toolkit $\mathcal{T}$, along with their input arguments and inter-dependencies. These subtasks are synthesized into a workflow represented as a Directed Acyclic Graph (DAG), $G$, where each node represents a task, and edges represent dependencies between them. E.g., a natural language question can be split into multiple tasks such as intent table detection, text2SQL, and image analysis as shown in Figure 2. The workflow reflects the execution sequence and dependencies that are necessary to answer the user’s query. The system also utilizes predefined prompts $\mathcal{P}$ to guide the reasoning process during task decomposition.
  2. (ii)
    Execution and Self-Debugging. The system executes the tasks according to the generated workflow by invoking the relevant expert models and tools from the toolkit $\mathcal{T}$. The system utilizes a state object $\mathcal{R}$, which stores intermediate results and interactions during the execution. The tasks are partitioned into independent batches $\mathcal{B}$ that can be executed in parallel, which is determined through a topological sort (TopologicalSort($G$)) of the DAG. For each batch, the system launches parallel executions of the assigned tasks. The tasks are executed using the expert models, and the outcomes are passed on to subsequent tasks that depend on them. Each expert model includes a self-debugging mechanism to detect and correct errors during execution. If an error persists, the system can provide feedback and retry the process, thereby enhancing the robustness of the execution.
  3. (iii)
    Decision Making. After the execution of the subtasks, M2EX inspects the intermediate results stored in $\mathcal{R}$ to determine whether they are sufficient to fulfill the user’s request. If the results are satisfactory, the system synthesizes them into the final answer $a$. However, if the results are insufficient or uncertain, the system triggers a re-planning process by invoking Replan($G,\mathcal{R},\mathcal{LLM},\mathcal{T}_{\text{meta}}$) to adjust the DAG and re-execute the tasks. This process repeats until the decision-making component is satisfied with the final result or a predefined maximum loop limit is reached.
In summary, M2EX uses an algorithmic approach where the system first decomposes the user query into subtasks, executes these tasks with error detection and correction mechanisms, and synthesizes the results into a final answer. The system is highly adaptive, with dynamic re-planning capabilities powered by the reasoning abilities of the LLM to ensure efficient task execution, debugging, and modification of the plan when needed. Our current M2EX implementation offers a range of features, including self-debugging, query re-planning, optimization, and explainability to better understand how a natural language question is decomposed into multiple sub-tasks. See details in Appendix D.
Complexity and convergence.
Planning inspects $|S|$ subtasks and calls $\Phi$ once, costing $\mathcal{O}(|S|\,C_{\text{LLM}})$ tokens ($C_{\text{LLM}}$ = context length processed by the language model). With unlimited workers, execution latency is $\mathcal{O}(\text{depth}(G))$; with $p$ workers it is bounded by $\text{depth}(G)$ as well. Re-planning only touches the affected sub-DAG, so its worst-case cost is strictly $\leq$ the first planning pass.
4 Experiments
In this section, we evaluate M2EX’s performance, focusing on the following research questions: (1) How well does M2EX tackle multi-modal natural language questions on three different datasets consisting of tabular data and images? (2) How does the system perform compared to state-of-the-art systems such as CAESURA (Urban and Binnig, 2024) and NeuralSQL (Bae et al., 2024) on underlying benchmark datasets? (3) What systematic errors can we observe?
4.1 Experimental Setup
Datasets For our experiments, we used three different datasets, namely datasets about artwork, basketball, as well as electronic health records. Due to hardware limitations, we reduced the dataset to 100 images and reports. Processing the full size in CAESURA can result in crashes due to out-of-memory issues.
  1. Dataset 1:
    ArtWork. This dataset was introduced by Urban and Binnig (2024) and contains information about paintings in tabular form as well as an image collection containing 100 images of the artworks, collected from Wikipedia. The tabular data contains metadata about paintings such as title, inception, movement, etc. as well as a reference to the respective paintings. A typical example question from this dataset is Plot the number of paintings depicting war for each century (see Figure 3 in the Appendix).